# Non-invertible duality defects in one, two, and three dimensions via gauging spatially modulated symmetry

### v1

@@ Abstract
Spatially modulated symmetries have emerged since the discovery of fractons, which characterize unconventional topological phases with mobility-constrained quasiparticle excitations. On the other hand, non-invertible duality defects have attracted substantial attention in communities of high energy and condensed matter physics due to their deep insight into quantum anomalies and exotic phases of matter. However, the connection between these exotic symmetries and defects has not been fully explored. In this paper, we construct concrete lattice models with non-invertible duality defects via gauging spatially modulated symmetries and investigate their exotic fusion rules. Specifically, we construct spin models with subsystem symmetries or dipole symmetries on one, two, and three-dimensional lattices. Gauging subsystem symmetries leads to non-invertible duality defects whose fusion rules involve 00-form subsystem charges in two dimensions and higher-form operators that correspond to “lineon” excitations (excitations which are mobile along one-dimensional line) in three dimensions. Gauging dipole symmetries leads to non-invertible duality defects with dipole algebras that describe a hierarchical structure between global and dipole charges. Notably, the hierarchical structure of the dual dipole charges is inverted compared with the original ones. Our work provides a unified and systematic analytical framework for constructing exotic duality defects by gauging relevant symmetries.
@@ 1 Introduction
Symmetry has been an important guiding principle in physics, allowing us to gain deeper insight of problems in various contexts. It imposes powerful constraints on a theory, from which one can predict many properties without explicit calculations. Recently, the concept of symmetries has been broadened, initiated by works [1, 2] which studied generalized symmetries, associated with extended objects, and we now have rich variety of symmetries which have been intensively investigated in different areas of research.
@@
In series of studies of fracton topological phases [3, 4, 5], unconventional topological phases admitting quasiparticle excitations with mobility constraints, there have been emerged new sets of symmetries, spatially modulated symmetries (a.k.a. exotic symmetries). One example of the spatially modulated symmetries is subsystem symmetry [6, 7, 5, 8], where a symmetry defect is topological in a particular direction, rather than entire space. A system with such a symmetry possesses a subextensive number of conserved charges, which is in line with the fact that the GSD becomes subextensive in fracton topological phases with subsystem symmetry, such as the X-cube model [5, 9, 10]. Further, subsystem symmetry gives rise to particle excitations with mobility constraints, such as “planons” (excitations which are mobile on a plane). Another typical example of the spatially modulated symmetries is multipole symmetry, a symmetry associated with conservation of multipoles, such as dipole [11, 12, 13, 14]. The multipole symmetry puts mobility constrains on excitations, giving rise to unusual GSD dependence on the system size in gapped phases, which was not observed in the case of the subsystem (see e.g., [15, 16, 17]). Moreover, the multipole symmetry yields new physical perspectives, allowing us to explore various research directions, such as field theories admitting multipole of fractional excitations [18, 19], the Hilbert space fragmentation in dynamically constrained systems [20, 21], and hydrodynamics [22, 23].
@@
Along a different stream of research, non-invertible symmetries (a.k.a. categorical symmetries) have attracted a plethora of attentions (see e.g., [24, 25, 26, 27, 28, 29, 30]). One familiar example of such a symmetry is the Kramers-Wannier duality defect in the Ising model [24]. Unified framework to deal with such symmetries has been developed by making use of fusion category theories [31, 32]. Although theoretical frameworks of the non-invertible symmetries in one dimension are well-established, elucidation of this type of symmetry in dimensions higher than one is an active area of research [33, 34, 35, 36, 37, 38].
@@
Among various kinds of symmetries, it is desirable to establish a theoretical scheme with broader scope, incorporating these symmetries in a unified fashion. Also, there are growing interests in explicit construction of non-invertible symmetry defects in lattice models. Spurred by these facts, in this paper, we explore the two types of symmetries, i.e., non-invertible symmetries and spatially modulated symmetries by constructing non-invertible duality defects obtained by gauging spatially modulated symmetry in one, two, and three dimensional lattice models. Our construction is motivated by recent progress of non-invertible defects in two dimensional systems comprised of double copies of spin models and gauging symmetry of the one of the degrees of freedom [39]. We find that non-invertible duality defects obtained by gauging subsystem or dipole symmetries exhibit rich structure of fusion rules. In the case of the subsystem symmetry, we argue that fusion rules of the duality defects are characterized by subsystem charges and higher form operators corresponding to the “lineon” excitations which are studied in the X-cube model [5], the simplest model of the fracton topological phase. Such fusion rules are demonstrated in (36) (61). Also, the duality defects obtained by gauging dipole symmetries, yield a new degree of freedom—dipole algebra, which describes how global and dipole charges are related via translational operators. We show that the hierarchical structure of such algebra plays a crucial role to understand fusion rules of the duality defects. Such fusion rules are summarized in (80), (112), (113), (135), and (136).
@@
Note that exploration of non-invertible duality defect via gauging spatially modulated symmetries has recently started (see e.g., [40, 41, 42, 43, 44]). Yet complete understanding of these defects, in particular defects in two and higher dimensions, remains elusive. Our work fills this gap and complies with diverse interests in various symmetries.
@@
The rest of this work is organized as follows. In Sec. 2, we go over how duality defect operator is obtained by the doubled copies of the spin model, the detail of which is crucial in the subsequent discussion. In Sec. 3, we construct non-invertible duality defect in two and three dimensions obtained by gauging subsystem symmetries. In Sec. 4, we turn our attention to the case of the dipole symmetries. After emphasizing the importance of the dipole algebra, we argue that fusion rules of non-invertible duality defects via gauging dipole symmetries involves such algebraic structure. In Sec. 5, we conclude our work with commenting on future perspectives. Technical issues are relegated into Appendix.
@@ 2 Conventional Ising chain
In this section, we briefly recap the argument presented in [39] to obtain systematically non-invertible duality defects. Specifically, we review non-invertible duality defects in double copies of the Ising chains, which also serves as the basic approach for our systematic constructions of non-invertible duality defects of spatially modulated symmetries in Sec. 3 and 4.
@@
To construct non-invertible duality defects, we gauge one of the global symmetries, say, $\tilde{Q}$. To do this, we introduce extended a Hilbert space, corresponding to gauge field, described by $\tilde{\tau}^{X}_{j+1/2}$, $\tilde{\tau}^{Z}_{j+1/2}$ located on each link. The Gauss law reads
@@
Defining new variables as
@@
Before gauging the global symmetry, the operator $S$ sends a local spin $X_{j}$ and $Z_{j}$ to $\tilde{X}_{j}$ and $\tilde{Z}_{j}$, respectively. These two are further mapped to $G^{X}_{j}$ and $\tau^{Z}_{j+1/2}$ after gauging, according to (9). Hence, one naively expects that the operator $S$ plays a role of the duality defects; operator $X_{j}$ is mapped to $G^{X}_{j}$ and $Z_{j}Z_{j+1}$ to $\tau^{Z}_{j+1/2}$, which can be seen in the typical duality transformation in the Ising model. However, one faces an issue; from the form of $S_{j}$ (3), a local operator $\tilde{Z}_{j}$ in $S_{j}$ does not commute with the global symmetry $\tilde{Q}$. Hence, after gauging, it should become non-local operator, involving the gauge fields $\tau^{Z}_{j+1/2}$. Based on the facts that the swap operator $S$ should behave as the duality defect, and that a local operator $\tilde{Z}_{j}$ inside the swap operator becomes nonlocal after gauging, we replace $\tilde{Z}_{j}$ with the Wilson operator $W_{j}$ consisting of string of the gauge fields:
@@
Note that $S_{j}~{}(1\leq j\leq L-1)$ is invertible but $S_{L}$ is non-invertible. Indeed, $S_{j}^{2}=1~{}(1\leq j\leq L-1)$ whereas $S_{L}^{2}=S_{L}$. Define product of the swap operator as $S=\prod_{j=1}^{L}S_{j}$, after some algebra one has
@@
The relations in (14) indicates that away from the boundary, $S$ indeed plays a role of the duality defect; yet the action of $S$ on spin at the boundary $(j=L)$ leads to non-local charges. To remedy this issue, one multiplies global charges with $S$ by introducing
@@
which acts as the desired duality defect:
$DX_{j}=G^{X}_{j}D,\quad DX_{j}=G^{X}_{j}D,\quad DZ_{j}Z_{j+1}=\tau^{Z}_{j+1/2}D,\quad D\tau^{Z}_{j+1/2}=Z_{j}Z_{j+1}D$ (16)
Furthermore, it is non-invertible, as is evident from the fusion rules of the defects
@@
The fusion rules (17) exhibits the non-invertible Rep$(D_{8})$ symmetry.
In the subsequent sections, we follow the approach discussed here for Ising models to construct non-invertible duality defects via gauging spatially modulated symmetries in lattice models. The strategy to establish such defects is proceeded in the following three steps:
@@
    Examine how the composite swap operator $S$ acts on spins and whether it behaves as the duality defect. The action of $S$ on some of spins may involve non-local charges, which is remedied by multiplying $S$ with global charges. [Shown in Eq. (14) to (17).]
As we demonstrate below, there are exotic structures in the duality defects obtained by gauging spatially modulated symmetries, such as 00-form and higher form subsystem charges and algebraic relations between global and dipole charges.
@@ 3 Subsystem symmetry
In this section, we construct non-invertible duality defects obtained by gauging subsystem symmetry in two and three dimensions.
@@ 3.1 Two dimensions
which is depicted in right of Fig. 1(a). Also, the four spins interaction defined on each plaquette is minimally coupled to the gauge field as  [^3] Regarding the original spin $\tilde{Z}_{\mathbf{r}}$ as a matter field, this procedure is practically associated with subsystem analog of the minimally coupling to the gauge fields [47].
@@
Analogous to the previous argument presented around (12) in the case of the double Ising chains, we expect the swap operator as the desired duality defect operator. However, we have issue with a local term, $\tilde{Z}_{\mathbf{r}}$ inside the operator $S_{\mathbf{r}}$ as it does not commute with the global symmetries (20). To fix this problem, we rewrite the term $\tilde{Z}_{\mathbf{r}}$ by a sting of gauge fields. Setting the coordinate lattice so that $\mathbf{r}_{0}=(0,0)$ as the reference point, we assume $\tilde{Z}_{\mathbf{r}}$ is transformed into product of $\tau^{Z}$, forming a rectangular, whose four corners are given by nodes at $\mathbf{r}=(\hat{x},\hat{y})$ and $\mathbf{r}_{0}$, $(\hat{x},0)$, and $(0,\hat{y})$. More explicitly, we replace $\tilde{Z}_{\mathbf{r}}$ with the following Wilson operator $W_{\mathbf{r}}$:
@@
The product of the swap operators (29) plays as the desired duality defect operator for spins, except the ones in vicinity of the end points of the lattice. Indeed, after some algebra, one finds that
@@
Relations (30) and (32) indicate that around the end points, the operator $S$ acts as a duality defect operator, up to charges of subsystem symmetries, such as $Q_{sub_{x},\hat{y}}$. To fix this issue, a proper duality operator is defined by multiplying projection operators with $S$, namely, we introduce the following:
@@
implying that the operator $D$ is the desired duality defect. Further, this duality defect is non-invertible; the fusion rules of the operators read
@@
Hence, we have constructed subsystem analog of the $\text{Rep}(D_{8})$ non-invertible duality defects (17). Here, the fusion rule of the duality defects yield subsystem charges.

### v2

# Noninvertible operators in one, two, and three dimensions via gauging spatially modulated symmetry

@@ Abstract
Spatially modulated symmetries have emerged since the discovery of fractons, which characterize unconventional topological phases with mobility-constrained quasiparticle excitations. On the other hand, noninvertible symmetry operators have attracted substantial attention in communities of high energy and condensed matter physics due to their deep insight into quantum anomalies and exotic phases of matter. However, the connection between these exotic and noninvertible symmetries has not been fully explored. In this paper, we construct concrete lattice models with noninvertible symmetry operators via gauging spatially modulated symmetries and investigate their exotic fusion rules. Specifically, we construct spin models with subsystem symmetries or dipole symmetries on one, two, and three-dimensional lattices. Gauging subsystem symmetries leads to noninvertible symmetry operators whose fusion rules involve 00-form subsystem charges in two dimensions and higher-form operators that correspond to “lineon” excitations (excitations which are mobile along one-dimensional line) in three dimensions. Gauging dipole symmetries leads to noninvertible symmetry operators with dipole algebras that describe a hierarchical structure between global and dipole charges. Notably, the hierarchical structure of the dual dipole charges is inverted compared with the original ones. Our work provides a unified and systematic analytical framework for constructing exotic symmetry operators by gauging relevant symmetries.
@@ 1 Introduction
Symmetry has been an important guiding principle in physics, allowing us to gain deeper insight of problems in various contexts—it imposes powerful constraints on a theory, from which one can predict many properties without explicit calculations. Recently, the concept of symmetries has been broadened, initiated by works [1, 2] on generalized symmetries associated with extended objects. With the advent of fractonic topological phases [3, 4, 5], new types of symmetries, namely, spatially modulated symmetries (a.k.a. exotic symmetries) emerge, where unconventional topological orders admit quasiparticle excitations with mobility constraints. In bosonic theories, the system has U(1) spatially modulated symmetry if the Hamiltonian is invariant under the transformation on the boson field $b_{\mathbf{r}}$ as $b_{\mathbf{r}}\to e^{i\theta(\mathbf{r})}b_{\mathbf{r}}$, where $\theta(\mathbf{r})$ depends on the lattice site $\mathbf{r}$. For instance, choosing $\theta(\mathbf{r})=a$ leads to the regular global U(1) symmetry; choosing $\theta(\mathbf{r})=\boldsymbol{b}\cdot\mathbf{r}$ leads to the dipole U(1) symmetry. Here $a$ and $\boldsymbol{b}$ are constants. The corresponding interactions with the global and dipole U(1) symmetries are $b^{\dagger}_{\mathbf{r}}b_{\mathbf{r}^{\prime}}$ and $b^{\dagger}_{\mathbf{r}}b_{\mathbf{r}+\hat{x}}b^{\dagger}_{\mathbf{r}+\hat{x}+\hat{y}}b_{\mathbf{r}+\hat{y}}$, respectively. Similarly, taking $\theta(\mathbf{r})$ as a step function leads to a subsystem symmetry. [^1] There is an important distinction between dipole and subsystem symmetries. In particular, space transformation act differently on unitary operators of these two symmetries. We denote $U(\Sigma)$ as a unitary operator that implements the unitary transformation dictated by $U$ on submanifold $\Sigma$. In the case of a subsystem symmetry, the unitary operator transforms as $SU(\Sigma)S^{\dagger}=U(\Sigma_{S}),$ (1) where $\Sigma_{S}$ represents another submaifold related by $\Sigma$ via spatial transformation $S$. On the other hand, in the case of a dipole symmetry, the unitary operator transforms as $SU(\Sigma)S^{\dagger}=U_{S}(\Sigma_{S})$ (2) In contrast to the transformation under a subsystem symmetry, $U_{S}$ is a different unitary transformation on the submanifold $\Sigma_{S}$. If one defines the modulated symmetry as a symmetry for which there is a group action of space symmetries on the internal symmetries such as the relation (2), dipole symmetries are modulated symmetries whereas subsystem symmetries are not. However, we regard both subsystem and dipole symmetries as modulated symmetries based on the fact that both have inhomogeneous symmetry transformations. An example for this inhomogeneous symmetry transformation is $\theta(\mathbf{r})$ in the case of the bosonic model explained above in the main text. This perspective is also adopted in the literature. See Ref. [6] for example. We thank the anonymous referee for pointing out this distinction.
@@
In this work, we focus on two examples of the spatially modulated symmetries – subsystem symmetry [7, 8, 5, 9], and multipole symmetries [10, 11, 12, 13]. In the former symmetry, a symmetry defect is topological in a particular direction, rather than entire space. Topological phases with subsystem symmetries possess a subextensive number of conserved charges, and the ground-state degeneracy (GSD) becomes subextensive in fracton topological phases with subsystem symmetry, such as the X-cube model [5, 14, 15]. Subsystem symmetries also lead to particle excitations with mobility constraints, such as “lineons” or “planons” (excitations which are mobile on a line or a plane, respectively). Multipole symmetries are associated with conservation of multipoles, such as dipole symmetries. These symmetries put mobility constrains on excitations, giving rise to unusual GSD dependence on the system size in gapped phases, which was not observed for subsystem symmetries (see e.g., [16, 17, 18]). Furthermore, multipole symmetries yield new insights in quantum field theories, allowing us to explore various research directions, such as multipolar fractional excitations [19, 20], Hilbert space fragmentations in dynamically constrained systems [21, 22], and hydrodynamics [23, 24].
@@
Along a different stream of research, noninvertible symmetries (a.k.a. categorical symmetries) have attracted a plethora of attentions (see e.g., [25, 26, 27, 28, 29, 30, 31]). One familiar example of such a symmetry is the Kramers-Wannier duality defect in the Ising model [25]. Unified framework to deal with such symmetries has been developed by making use of fusion category theories [32, 33]. Although theoretical frameworks of the noninvertible symmetries in one dimension are well-established, elucidation of this type of symmetries in higher dimensions is still an active area of research [34, 35, 36, 37, 38, 39, 40, 41].
@@
Among various kinds of symmetries, it is desirable to establish a theoretical framework with broader scope, incorporating these symmetries in a unified fashion. On the other hand, there are growing interests in explicit constructions of noninvertible symmetry operators in lattice models. Spurred by these facts, in this paper, we explore intertwining of the two types of symmetries, i.e., noninvertible symmetries and spatially modulated symmetries by constructing exactly solvable lattice models of noninvertible operators by gauging spatially modulated symmetries in one, two, and three dimensions. Our construction is motivated by recent progress of noninvertible operators in two dimensional systems comprised of double copies of spin models [42]. We find that noninvertible operators obtained by gauging subsystem or dipole symmetries exhibit rich structures in fusion rules. For subsystem symmetries, we show that fusion rules of the operators are characterized by subsystem charges and higher form operators corresponding to the “lineon” excitations which are studied in the X-cube model [5], the simplest model of the fracton topological phase. Such fusion rules are demonstrated in (41) (67). For dipole symmetries, the operators exhibit a new type of dipole algebras, which describes how global and dipole charges are related via translational operators. We show that the hierarchical structure of such algebra plays a crucial role to understand fusion rules of the operators. Such fusion rules are summarized in (87), (121), (122), (144), and (145).
@@
Note that explorations of noninvertible operators via gauging spatially modulated symmetries have recently started (see e.g., [43, 44, 45, 46, 47, 48, 49, 50]). Yet complete understanding of these operators, in particular operators in two and higher dimensions, remains elusive. Our work fills this gap and complies with diverse interests in various symmetries.
@@
The rest of this work is organized as follows. In Sec. 2, we go over how noninvertible operator is obtained by the doubled copies of the spin model, the detail of which is crucial in the subsequent discussion. In Sec. 3, we construct noninvertible operator in two and three dimensions obtained by gauging subsystem symmetries. In Sec. 4, we turn our attention to the case of the dipole symmetries. After emphasizing the importance of the dipole algebra, we argue that fusion rules of noninvertible operators via gauging dipole symmetries involves such algebraic structure. In Sec. 5, we conclude our work with commenting on future perspectives. Technical issues are relegated into Appendix.
@@ 2 Conventional Ising chain
In this section, we briefly recap the argument presented in [42] to obtain systematically noninvertible operators. Specifically, we review noninvertible operators in double copies of the Ising chains, which also serves as the basic approach for our systematic constructions of noninvertible operators of spatially modulated symmetries in Sec. 3 and 4.
@@
To construct noninvertible operators, we gauge one of the global symmetries, say, $\tilde{Q}$. To do this, we introduce extended a Hilbert space, corresponding to gauge field, described by $\tilde{\tau}^{X}_{j+1/2}$, $\tilde{\tau}^{Z}_{j+1/2}$ located on each link. The Gauss law reads [^2] Alternatively, one performs a unitary transformation (see e.g.,[52]) that simplifies the Gauss law term to a local single $\mathbb{Z}_{2}$ operator, say, $\zeta^{X}_{j}$ and set $\zeta^{X}_{j}=1$.
@@
Defining new gauge invariant variables as
@@
Before gauging the global symmetry, the operator $S$ sends a local spin $X_{j}$ and $Z_{j}$ to $\tilde{X}_{j}$ and $\tilde{Z}_{j}$, respectively. These two are further mapped to $G^{X}_{j}$ and $\tau^{Z}_{j+1/2}$ after gauging, according to (11). Hence, one naively expects that the operator $S$ plays a role of the operators; operator $X_{j}$ is mapped to $G^{X}_{j}$ and $Z_{j}Z_{j+1}$ to $\tau^{Z}_{j+1/2}$, which can be seen in the typical duality transformation in the Ising model. However, one faces an issue; from the form of $S_{j}$ (5), a local operator $\tilde{Z}_{j}$ in $S_{j}$ does not commute with the global symmetry $\tilde{Q}$. Hence, after gauging, it should become non-local operator, involving the gauge fields $\tau^{Z}_{j+1/2}$. Based on the facts that the swap operator $S$ should behave as the operator, and that a local operator $\tilde{Z}_{j}$ inside the swap operator becomes nonlocal after gauging, we replace $\tilde{Z}_{j}$ with the Wilson operator $W_{j}$ consisting of string of the gauge fields:
@@
Note that $S_{j}~{}(1\leq j\leq L-1)$ is invertible but $S_{L}$ is noninvertible. Indeed, $S_{j}^{2}=1~{}(1\leq j\leq L-1)$ whereas $S_{L}^{2}=S_{L}$. Also, $\{S_{j}|1\leq j\leq L-1\}$ commutes with one another whereas $S_{L}$ does not with $S_{j}\;(1\leq j\leq L-1)$. Define product of the swap operator as $S=\prod_{j=1}^{L}S_{j}$, after some algebra one has
@@
Each relation in (16) can be verified by a simple algebra. For instance, given $SX_{i}=\tau^{x}_{i-1/2}\tau^{x}_{i+1/2}$ for $i=1,\dots,L-1$, we have
$\tau^{x}_{L-1/2}\tau^{x}_{1/2}S$ $=(\tau^{x}_{L-1/2}\tau^{x}_{L-3/2})\cdot(\tau^{x}_{L-3/2}\tau^{x}_{L-5/2})\cdots(\tau^{x}_{3/2}\tau^{x}_{1/2})S$
$=SX_{L-1}X_{L-2}\cdots X_{1}=SX_{L}Q,$ (17)
which is essentially
$SX_{L}$ $=G^{X}_{L}SQ.$ (18)
The relations in (16) indicates that away from the boundary, $S$ indeed plays a role of the mapping corresponding to gauging (11); yet the action of $S$ on spin at the boundary $(j=L)$ leads to non-local charges. To remedy this issue, one multiplies global charges with $S$ by introducing
@@
which acts as the desired swap operator:
$DX_{j}=G^{X}_{j}D,\quad DG^{X}_{j}=X_{j}D,\quad DZ_{j}Z_{j+1}=\tau^{Z}_{j+1/2}D,\quad D\tau^{Z}_{j+1/2}=Z_{j}Z_{j+1}D$ (20)
Furthermore, it is noninvertible, as is evident from the fusion rules of the operators
@@
The fusion rules (21) exhibits the noninvertible Rep$(D_{8})$ symmetry.
In the subsequent sections, we follow the approach discussed here for Ising models to construct noninvertible operators via gauging spatially modulated symmetries in lattice models. The strategy to establish such operators is proceeded in the following three steps:
@@
    Examine how the composite swap operator $S$ acts on spins and whether it behaves as the mapping corresponding to gauging. The action of $S$ on some of spins may involve non-local charges, which is remedied by multiplying $S$ with global charges. [Shown in Eq. (16) to (21).]
As we demonstrate below, there are exotic structures in the operators obtained by gauging spatially modulated symmetries, such as 00-form and higher form subsystem charges and algebraic relations between global and dipole charges.
@@ 3 Subsystem symmetry
In this section, we construct noninvertible operators obtained by gauging subsystem symmetry in two and three dimensions.
@@ 3.1 Two dimensions
which is depicted in right of Fig. 1(a). Intuition behind the form of the Gauss law (27) is that one decomposes global symmetries into local segments with introducing the extended Hilbert spaces (which are associated with the degrees of freedom of the gauge fields), corresponding to the fact that gauging is a procedure to promote global symmetries to local ones. Indeed, defining
$\mathcal{G}_{sub,\mathbf{r}}\vcentcolon=\widehat{\tau}^{X}_{\mathbf{r}+\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}}\widehat{\tau}^{X}_{\mathbf{r}+\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}}\widehat{\tau}^{X}_{\mathbf{r}-\frac{\mathbf{e}_{x}}{2}+\frac{\mathbf{e}_{y}}{2}}\widehat{\tau}^{X}_{\mathbf{r}-\frac{\mathbf{e}_{x}}{2}-\frac{\mathbf{e}_{y}}{2}}\tilde{X}_{\mathbf{r}}$
one has
$\tilde{Q}_{sub_{y},\hat{x}}=\prod_{\hat{y}=1}^{L_{y}}\mathcal{G}_{sub,\mathbf{r}}\;(1\leq\hat{x}\leq L_{x}),\quad\tilde{Q}_{sub_{x},\hat{y}}=\prod_{\hat{x}=1}^{L_{x}}\mathcal{G}_{sub,\mathbf{r}}\;(1\leq\hat{y}\leq L_{y}).$ (28)
The Gauss law is imposed by setting $\mathcal{G}_{sub,\mathbf{r}}=1\quad\forall\mathbf{r}$. In short, the Gauss law in Eq. (27) is derived from the standard gauging procedure, i.e. promoting the global invariance into the local invariance and then minimally coupling the system to background gauge fields which we turn to momentarily.
The four spins interaction defined on each plaquette is minimally coupled to the gauge field as  [^5] Regarding the original spin $\tilde{Z}_{\mathbf{r}}$ as a matter field, this procedure is practically associated with subsystem analog of the minimally coupling to the gauge fields [54].
@@
Analogous to the previous argument presented around (14) in the case of the double Ising chains, we expect the swap operator as the desired noninvertible operator. However, we have issue with a local term, $\tilde{Z}_{\mathbf{r}}$ inside the operator $S_{\mathbf{r}}$ as it does not commute with the global symmetries (24). To fix this problem, we rewrite the term $\tilde{Z}_{\mathbf{r}}$ by a string of gauge fields. Setting the coordinate lattice so that $\mathbf{r}_{0}=(0,0)$ as the reference point, we assume $\tilde{Z}_{\mathbf{r}}$ is transformed into product of $\tau^{Z}$, forming a rectangular, whose four corners are given by nodes at $\mathbf{r}=(\hat{x},\hat{y})$ and $\mathbf{r}_{0}$, $(\hat{x},0)$, and $(0,\hat{y})$. More explicitly, we replace $\tilde{Z}_{\mathbf{r}}$ with the following Wilson operator $W_{\mathbf{r}}$:
@@
The product of the swap operators (34) plays as the desired noninvertible operator for spins, except the ones in vicinity of the end points of the lattice. Indeed, after some algebra, one finds that
@@
Relations (35) and (37) indicate that around the end points, the operator $S$ acts as a mapping corresponding to gauging, up to charges of subsystem symmetries, such as $Q_{sub_{x},\hat{y}}$. To fix this issue, a proper operator is defined by multiplying projection operators with $S$, namely, we introduce the following:
@@
implying that the operator $D$ is the desired swap operator. Further, this operator is noninvertible; the fusion rules of the operators read
@@
Hence, we have constructed subsystem analog of the $\text{Rep}(D_{8})$ noninvertible operators (21). Here, the fusion rule of the operators yield subsystem charges.